Technology Seminar, Math 598, Thursdays, 11am-noon

Date  Speaker(s)  Topic(s)
June 22  Felix Huang  Complex Dynamics
July 6  Andreas Wiegmann Converting dvi to html
July 13  Paul Moorehead  Zipper
July 20  Don Bovee  Web Stuff + Questions and Answers
July 27  David Ingerman  Brownian Motion
August 3  Don McConnell  Clickable Images
August 10  Carlos Tolmasky  Inmersed Interface Method and Landmines
August 17  Tim Roberts  Home computing
Last revised: August 3, 1995